# Zhang, Liu, and Zhao

### v3

@@ 6.3 Nonconvex Newsvendor Problem
Newsvendor Problem with a Nonconvex Objective
b. Nonconvex Constraint. We next consider the 2-product Newsvendor problem given in Section 6.2 with parameters $(c_{b1},c_{h1},c_{b2},c_{h2})=(7,7,3,3)$ and the replacement of the convex constraint by $C(z_{1})+C(z_{2})\leq C_{0}$. We set $C_{0}=50$ to keep the same active probability for the constraint. Since the projection of infeasible decisions to the nonconvex feasible region may not yield meaningful solutions, we only report the average test cost of feasible decisions in Figure 1o, with a careful hyperparameter selection to ensure high feasible frequencies.
(m) Basic Results
(n) Sparse Demand Model
(o) Dense Demand Model
Newsvendor Problem with a Nonconvex Constraint In (b) and (c) with $p=50$, RKHS-DR(MM) takes more than 500 seconds on average, while PADR(2,2) takes less than 200 seconds.
For the problem with the basic demand model (12), PADR(2,2) performs comparably with RKHS-DR(MM), with a notable gap to SIMOPT. Figures 1n and 1o indicate the superiority of PADR(2,2) over the other two methods under both sparse and dense demand models with varying feature dimensions. It’s also worth noting that the simpler PADR(1,0) (equivalent to LDR), with poor performance in general, performs comparably well with the others in the high-dimensional case. It is more difficult to obtain feasible solutions for cases with sparse models than those with dense models; moreover, with sparse models, PADR(1,0) is more likely to produce feasible solutions than the others.
Finally, we also conduct experiments on constrained Newsvendor problems in Sections 6.2 and 6.3 with multiple $C_{0}$ values and Newsvendor cost parameters. The numerical results exhibit similar patterns and thus are omitted.
@@ 7 Conclusion
We propose the PADR-based ERM method together with the enhanced SMM algorithm to solve a broad class of SP problems with covariate information. For our PADR-based ERM framework, we provide the approximation bound of the PA function class and further establish the nonasymptotic consistency result for unconstrained problems and the asymptotic consistency result for deterministic-constrained ones. To solve the ERM problem with a composite DC structure, we develop the enhanced SMM algorithm by combining enhanced $\varepsilon$-active surrogation with sequential sampling. We provide the asymptotic convergence of the enhanced SMM algorithm to composite strong d-stationarity, and the nonasymptotic convergence to d-stationarity via the error-bound analysis. Numerical experiments indicate the superiority of our method over PO methods and prominent DR-based methods in the Newsvendor problem regarding the test cost, computational time, and robustness to feature dimensions and nonlinearity of the underlying dependency. In conclusion, the PADR is a favorable hypothesis class for DR-based ERM methods, supported by theoretical consistency, algorithmic convergence, and computational efficiency for solving SP with covariate information. Extensions of the PADR-based ERM framework for more complicated SP problems are deserved to be explored in the future, such as problems with chance constraints or endogenous random variables, which necessitate a further investigation of model construction, consistency analysis, and data-driven algorithmic development for problems with more complex composite structures.
@@ 8 Proofs of Theoretical Consistency Results in Section 2
The idea of the proof is to construct an interpolation PA function to a set of points related to the target function $f_{0}$ and a cover of the set $\mathcal{X}$. The structure of the interpolation function is motivated by Siahkamari et al. (2020) but with a different parameter, which ensures the Lipschitz continuity of the function with modulus independent of the approximation accuracy.
Proof 8.1
@@
Given $f_{0}\in\mathcal{F}_{L_{0},M_{0}}$, by McShane (1934, Theorem 1), we can extend $f_{0}$ to an $L_{0}$-Lipschitz function $\bar{f}_{0}:\bar{\mathcal{X}}\rightarrow\mathbb{R}$ satisfying $\left\lVert\bar{f}_{0}\right\rVert_{\infty}\leq M_{0}$ and $\bar{f}_{0}(x)=f_{0}(x)$ for any $x\in\mathcal{X}$. Next, we construct a piecewise affine interpolation function to a set of points $\{(\widehat{x},\bar{f}_{0}(\widehat{x}))\mid\widehat{x}\in\widehat{\mathcal{C}}_{2}(\epsilon;\bar{\mathcal{X}})\}$:
@@
Let $\Phi$ be a class of functions mapping from $\mathcal{U}$ to $\mathbb{R}$ and $S_{N}=\{u^{s}\}_{s=1}^{N}\subseteq\mathcal{U}$ be a sample set of size $N$. Then the empirical Rademacher complexity of $\Phi$ with respect to $S_{N}$ is defined as
^$\widehat{\mathcal{R}}_{S_{N}}(\Phi)\coloneqq\mathbb{E}_{\boldsymbol{\sigma}}\left[\sup_{\phi\in\Phi}\frac{1}{N}\sum_{s=1}^{N}\sigma_{s}\phi(u^{s})\right],$
@@
$\mathcal{R}_{N}(\Phi)\coloneqq\mathbb{E}_{S_{N}}[\widehat{\mathcal{R}}_{S_{N}}(\Phi)].$
It captures the richness of the class $\Phi$ by measuring the degree to which a hypothesis set can fit random noise. The following lemma quantifies the Rademacher complexity of a class of parametric functions, which is generalized from Ermoliev and Norkin (2013, Lemma B.2) with minor modifications made to accommodate the developments in this paper.
@@
Proof 8.4
Proof of Lemma 8.3 Given a sample set $S_{N}=\{u^{s}\}_{s=1}^{N}$ of size $N$, following the proof of Ermoliev and Norkin (2013, Lemma B.2), we derive that for any $\epsilon\in(0,2\mu]$,
^$\widehat{\mathcal{R}}_{S_{N}}(\widehat{\Phi})\leq L_{\widehat{\Phi}}\sqrt{q}\epsilon+M_{\widehat{\Phi}}\sqrt{\frac{2q\log(2\mu/\epsilon)}{N}}.$
By setting $\epsilon=2\mu/N$ and taking expectation over $S_{N}$, we obtain the desired upper bound.∎
@@
Lemma 8.5 (Mohri et al. (2018), Theorem 3.3)
Let $\Phi$ be a function class mapping from $\mathcal{U}$ to $[-M,M]$. Then for any $\delta>0$, with probability at least $1-\delta$ over the draw of an independently and identically distributed (i.i.d.) sample $S_{N}$ of size $N$, the following holds true for all $\phi\in\Phi$:
@@
Proof 8.6
Proof of Proposition 2.2. We first quantify the Rademacher complexity of the composite function class $\Phi_{{\mathcal{H}}_{\mu}^{K}}\coloneqq\left\{\phi:\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}\mid\phi(x,y)=\varphi({f(x)};{y}),\,f\in{\mathcal{H}}_{\mu}^{K}\right\}$. Notice that any function in $\Phi_{{\mathcal{H}}_{\mu}^{K}}$ is parameterized by some $\theta\in[-\mu,\mu]^{2dK(p+1)}$. For any $(x,y)\in\mathcal{X}\times\mathcal{Y}$ and $\theta_{1},\theta_{2}\in[-\mu,\mu]^{2dK(p+1)}$,
$\left\lvert\varphi(f(x;\theta_{1});y)-\varphi(f(x;\theta_{2});y)\right\rvert\leq L_{\varphi}\left\lVert f(x;\theta_{1})-f(x;\theta_{2})\right\rVert\leq 2L_{\varphi}\sqrt{p\bar{X}^{2}+1}\left\lVert\theta_{1}-\theta_{2}\right\rVert,$
@@
$\left\lvert\varphi(f(x);y)\right\rvert\leq\left\lvert\varphi(z_{0};y)\right\rvert+\left\lvert\varphi(f(x);y)-\varphi(z_{0};y)\right\rvert\leq C_{\varphi}+L_{\varphi}\left\lVert f(x)-z_{0}\right\rVert\leq M_{\mu}.$ (14)
@@
$\max_{f\in{\mathcal{H}}_{\mu}^{K}}\left\lvert R(f)-\widehat{R}_{S}(f)\right\rvert\leq 2\mathcal{R}_{n}(\Phi_{{\mathcal{H}}_{\mu}^{K}})+\sqrt{2}M_{\mu}\sqrt{\frac{\log(2/\delta)}{n}}.$ (16)
@@
$=\mathbb{E}_{X,Y}\left[\varphi({f_{app}(X)};{Y})-\varphi({f^{*}(X)};{Y})\right]$
@@
$R(f^{*}_{S,{\mathcal{H}}_{\mu}^{K}})-R(f^{*}_{{\mathcal{H}}_{\mu}^{K}})$ $=R(f^{*}_{S,{\mathcal{H}}_{\mu}^{K}})-\widehat{R}_{S}(f^{*}_{S,{\mathcal{H}}_{\mu}^{K}})+\widehat{R}_{S}(f^{*}_{S,{\mathcal{H}}_{\mu}^{K}})-R(f^{*}_{{\mathcal{H}}_{\mu}^{K}})$
$\leq R(f^{*}_{S,{\mathcal{H}}_{\mu}^{K}})-\widehat{R}_{S}(f^{*}_{S,{\mathcal{H}}_{\mu}^{K}})+\widehat{R}_{S}(f^{*}_{{\mathcal{H}}_{\mu}^{K}})-R(f^{*}_{{\mathcal{H}}_{\mu}^{K}})$
$\leq 2\left(\max_{f\in{\mathcal{H}}_{\mu}^{K}}\,|R(f)-\widehat{R}_{S}(f)|\right).$
@@
$R(f^{*}_{S,{\mathcal{H}}_{\mu}^{K}})-R(f^{*}_{{\mathcal{H}}_{\mu}^{K}})\leq\frac{16\mu L_{\varphi}\sqrt{2dK(p+1)(p\bar{X}^{2}+1)}}{n}+4M_{\mu}\sqrt{\frac{2dK(p+1)\log n}{n}}+2\sqrt{2}M_{\mu}\sqrt{\frac{\log(2/\delta)}{n}}.$
@@
9 Surrogation Construction of $F(\cdot;\xi)$ for Case 1 to Case 3
In this section, we first illustrate the construction of upper surrogation functions of $F(\cdot;\xi)\coloneqq\varphi(f(\cdot;\xi);\xi)$ for Case 1 to Case 3, which are directionally differentiable and satisfy the first three properties in Assumption 1. Then we show in Claim 1 that $\widehat{F}^{\varepsilon}_{\min}(\theta;\theta^{\prime},\xi)\coloneqq\min_{I\in\mathcal{I}^{\varepsilon}(\theta^{\prime})}\widehat{F}(\theta;\theta^{\prime},\xi,I)$ with $\varepsilon\geq 0$ satisfies Condition (P5) for the three cases, which also validates Condition (P4) since $\widehat{F}^{\varepsilon}_{\min}(\cdot;\theta^{\prime},\xi)$ by construction is directionally differentiable and satisfies (P2).
Case 1:
since $f(\cdot;\xi)$ and $\varphi(\cdot;\xi)$ are both piecewise affine, the composite function $F(\cdot;\xi)=\varphi(f(\cdot;\xi);\xi)$ is also piecewise affine. Thus, following the surrogation construction of $f(\cdot;\xi)$ with the difference-of-max-affine representation in Section 3.1, we can construct the upper surrogation function of $F(\cdot;\xi)$, which satisfies the first three properties in Assumption 1. Specifically, the first two properties clearly hold, and the bi-variate continuity of $\widehat{F}(\cdot;\cdot,\xi,I)$ is justified since given a fixed index mapping $I$, the surrogate of a PA function is independent of the reference point $\theta^{\prime}$.
Case 2:
by assumption, $\varphi(\cdot;\xi)$ can be written as $\varphi(\cdot;\xi)=\varphi^{\uparrow}(\cdot;\xi)+\varphi^{\downarrow}(\cdot;\xi)$, where $\varphi^{\uparrow}(\cdot;\xi)$ and $\varphi^{\downarrow}(\cdot;\xi)$ are nondecreasing and nonincreasing convex functions respectively, that is, for any $\theta_{1},\theta_{2}\in\Theta$ satisfying $\theta_{1}\leq\theta_{2}$ componentwisely, we have $\varphi^{\uparrow}(\theta_{2};\xi)\geq\varphi^{\uparrow}(\theta_{1};\xi)$ and $\varphi^{\downarrow}(\theta_{2};\xi)\leq\varphi^{\downarrow}(\theta_{1};\xi)$. With $I=(I_{1},I_{2})\in\mathcal{I}^{\varepsilon}(\theta^{\prime})$, the surrogation function of $F(\cdot;\xi)$ at $\theta^{\prime}$ is constructed as
$\widehat{F}(\theta;\theta^{\prime},\xi,I)=\varphi^{\uparrow}(\widehat{f}(\theta;\xi,I_{2});\xi)+\varphi^{\downarrow}(\widecheck{f}(\theta;\xi,I_{1});\xi),$
where $\widehat{f}(\theta;\xi,I_{2})$ and $\widecheck{f}(\theta;\xi,I_{1})$ are the upper and lower surrogation functions of $f(\theta,\xi)$ respectively. By monotonicity of $\varphi^{\uparrow}$ and $\varphi^{\downarrow}$, $\widehat{F}(\theta;\theta^{\prime},\xi,I)\geq\varphi^{\uparrow}(f(\theta;\xi);\xi)+\varphi^{\downarrow}(f(\theta;\xi);\xi)=F(\theta;\xi)$ for any $\theta\in\Theta$, and thus the majorizing condition (P2) holds. Moreover, $\widehat{F}(\cdot;\theta^{\prime},\xi,I)$ satisfies the convex condition (P3) since a nondecreasing convex function composite with a convex function and a nonincreasing convex function composite with a concave function are both convex. The touching condition (P1) clearly holds when $\varepsilon=0$. With a fixed index mapping $I$, $\widehat{F}(\cdot;\theta^{\prime},\xi,I)$ in this case is also independent of $\theta^{\prime}$, validating the bi-variate continuity of $\widehat{F}(\cdot;\cdot,\xi,I)$.
Case 3:
since $S$ is finite, for any $\xi\in S$, $\varphi(\cdot;\xi)$ is Lipschitz smooth with uniform modulus denoted by $L_{\nabla\varphi}$, and $f(\cdot;\xi)$ is Lipschitz continuous with uniform modulus $L_{f}\coloneqq 2\sqrt{\max_{s\in[n]}\left\lVert x^{s}\right\rVert^{2}+1}$. Then the surrogation function of $F(\cdot;\xi)$ at $\theta^{\prime}$ with $I=(I_{1},I_{2})\in\mathcal{I}^{\varepsilon}(\theta^{\prime})$ is constructed by surrogating the inner PA function according to componentwise signs of the gradient $\nabla\varphi(f(\theta^{\prime};\xi);\xi)$ as follows,
$\widehat{F}(\theta;\theta^{\prime},\xi,I)$ $=\max\{\nabla{\varphi}(f(\theta^{\prime};\xi);\xi),0\}^{\top}\widehat{f}(\theta;\xi,I_{2})-\max\{-\nabla{\varphi}(f(\theta^{\prime};\xi);\xi),0\}^{\top}\widecheck{f}(\theta;\xi,I_{1})$
$\phantom{=,}-\nabla{\varphi}(f(\theta^{\prime};\xi);\xi)^{\top}f(\theta^{\prime};\xi)+\varphi(f(\theta^{\prime};\xi);\xi)+\frac{1}{2}L_{\nabla\varphi}L_{f}^{2}\left\lVert\theta-\theta^{\prime}\right\rVert^{2}.$
The convex condition (P3) holds evidently; the touching condition (P1) also holds when $\varepsilon=0$. The surrogate $\widehat{F}(\cdot;\cdot,\xi,I)$ is continuous on $\Theta\times\Theta$ since $\nabla\varphi(\cdot;\xi)$ is continuous. Based on the quadratic upper bound of a Lipschitz smooth function, the majorizing condition (P2) is verified as follows. Since $\varphi(\cdot;\xi)$ is $L_{\nabla\varphi}$-smooth, we have for any $z_{1},z_{2}\in\mathbb{R}^{d}$ that
$\varphi(z_{1};\xi)\leq\varphi(z_{2};\xi)+\nabla\varphi(z_{2};\xi)^{\top}(z_{1}-z_{2})+\frac{L_{\nabla\varphi}}{2}\left\lVert z_{1}-z_{2}\right\rVert^{2}.$
By setting $z_{1}=f(\theta;\xi)$ and $z_{2}=f(\theta^{\prime};\xi)$, we obtain that for any $I=(I_{1},I_{2})\in\mathcal{I}^{\varepsilon}(\theta^{\prime})$,
$F(\theta;\xi)$ $\leq F(\theta^{\prime};\xi)+\nabla\varphi(f(\theta^{\prime};\xi);\xi)^{\top}(f(\theta;\xi)-f(\theta^{\prime};\xi))+\frac{L_{\nabla\varphi}}{2}\left\lVert f(\theta;\xi)-f(\theta^{\prime};\xi)\right\rVert^{2}$
$\leq F(\theta^{\prime};\xi)+\max\{\nabla{\varphi}(f(\theta^{\prime};\xi);\xi),0\}^{\top}\widehat{f}(\theta;\xi,I_{2})$
$\phantom{\leq,}-\max\{-\nabla{\varphi}(f(\theta^{\prime};\xi);\xi),0\}^{\top}\widecheck{f}(\theta;\xi,I_{1})-\nabla\varphi(f(\theta^{\prime};\xi);\xi)^{\top}f(\theta^{\prime};\xi)+\frac{L_{\nabla\varphi}L_{f}^{2}}{2}\left\lVert\theta-\theta^{\prime}\right\rVert^{2}$
$=\widehat{F}(\theta;\theta^{\prime},\xi,I).$
Claim 1
Let $\Theta$ be a compact set. Given $\varepsilon\geq 0$ and $\theta^{\prime}\in\Theta$, with the surrogation construction for Case 1 to Case 3, $\widehat{F}^{\varepsilon}_{\min}(\cdot;\theta^{\prime},\xi)$ satisfies Condition (P5) for any $\xi\in S$.
Proof 9.1
Proof of Claim 1. Since $\widehat{F}^{\varepsilon}_{\min}(\cdot;\theta^{\prime},\xi)$ also satisfies the majorizing condition (P2), by the compactness of $\Theta$, it suffices to show that

### v4

@@ 4.1 Newsvendor problem: simulated data
4.2 Newsvendor problem with nonconvex cost: real-world data
In this section, we conduct experiments on a real-world dataset, which consists of order records for a specific model of mobile phone from a Chinese electronic product supplier. The order record, spanning from September 01, 2021 to November 31, 2023, contains order quantities and time stamps from retailers. In order to create a production plan, the supplier needs to determine the packing quantity in each day to fulfill the order requests.
We first organize the individual order records into daily order quantities. The aggregated daily order quantities have been rescaled for confidentiality purposes and also normalized for pretreatment. The boxplots of daily orders in Figure 1n show that order quantities are larger on weekdays and during the second quarter of each month, which aligns with retailers’ ordering pattern.
[…]
(l) Weekdays
(m) Month days
(n) Month quarters
Real-world dataset: boxplots of orders by time period
We set observations from the first year as the training set and the remaining data as the test set. Here we consider three types of features, namely weekday, month quarter, and order quantities in past days. The first two types of features are converted into dummy variables (7 for weekdays and 4 for month quarters) and set as basic features. For the third type of features, we incorporate order quantities in past days of different lengths as different instances in conducting the experiments. To determine the packing quantity in each day to fulfill the order requests, we formulate a Newsvendor problem with features listed above and report the numerical results in Table 4.2. It shows that PADR and forest-based prescriptive methods work well. Overall, the PADR method performs slightly better in most instances with features of different length.
[…]
Test costs for newsvendor on a real-world dataset Features $p$ DR methods Prescriptive methods PO methods PADR RKHS-DR kNN cart rf sof PO-PA basic features 11 55.95 56.10 56.06 56.00 56.00 56.00 79.02 basic features + order quantities in past 1 day 13 42.67 51.53 50.25 61.72 49.54 48.70 68.18 basic features + order quantities in past 3 days 15 41.65 55.59 48.25 54.64 41.74 42.19 59.73 basic features + order quantities in past 5 days 17 41.40 65.17 49.59 56.74 40.90 41.83 52.39 basic features + order quantities in past 7 days 19 40.93 70.60 49.02 65.62 43.62 43.51 54.52 basic features + order quantities in past 14 days 26 41.03 78.37 46.81 58.39 42.45 42.48 52.10 basic features + order quantities in past 21 days 33 42.43 80.24 46.61 59.45 41.59 41.33 53.44 The lowest three costs in each row are highlighted in bold and color, with the lowest one being underlined and colored red.
We next consider a practical setting where the supplier incurs an additional cost associated with packing quantities in actual production operations, in order to balance the daily inventory cost and the packing cost. The cost is approximated by the following nonconvex cost function:
$C(z)=\min\{z,0.6z+0.8,0.4z+15.6\},$ (18)
Therefore, we consider a Newsvendor problem with an additional concave cost $C(z)$ in the objective. Since the concave cost introduces nonconvexity into the problem. we use the deterministic MM algorithm to solve the ERM problem for learning RKHS-based decision rule, referred to as RKHS-DR(MM). For prescriptive methods, we solve one-dimensional nonconvex problems by line search. We do not implement sof from Kallus and Mao (2023) in this nonconvex Newsvendor problem, as it requires a new development of second-order perturbation analysis of stochastic optimization, which is beyond the scope of this paper.
[…]
Test costs for nonconvex newsvendor on a real-world dataset
Features $p$ DR methods Prescriptive methods PO methods
PADR RKHS-DR(MM) kNN cart rf PO-PA
basic features 11 74.28 81.89 84.52 81.56 81.56 90.48
basic features + order quantities in past 1 day 13 64.75 72.11 77.86 86.58 73.78 77.89
basic features + order quantities in past 3 days 15 64.10 68.29 75.58 78.31 64.99 73.61
basic features + order quantities in past 5 days 17 63.20 64.72 76.81 79.54 65.74 71.56
basic features + order quantities in past 7 days 19 62.01 65.35 75.38 85.95 68.35 69.97
basic features + order quantities in past 14 days 26 64.86 65.15 73.81 75.52 64.71 70.80
basic features + order quantities in past 21 days 33 65.65 65.15 73.81 75.52 65.73 71.02
The lowest three costs in each row are highlighted in bold and color, with the lowest one being underlined and colored red.
[…]
4.3 Product placement problem
We consider a product placement problem on a graph $\mathcal{G}=(\mathcal{N},\mathcal{A})$ with the aim to determine the quantities of products $z\in\mathbb{R}^{\left\lvert\mathcal{N}\right\rvert}$ to place at different nodes. After placing the products, demands $y\in\mathbb{R}^{\left\lvert\mathcal{N}\right\rvert}$ are observed at nodes $\mathcal{N}$, and then products are shipped across arcs $\mathcal{A}$ to fulfill these demands. We denote the node-arc matrix of $\mathcal{G}$ by $W$, the product placement cost by $c\in\mathbb{R}^{\left\lvert\mathcal{N}\right\rvert}$, the shipment cost along edges by $g\in\mathbb{R}^{\left\lvert\mathcal{A}\right\rvert}$, and the penalty for unmet demand by $b\in\mathbb{R}^{\left\lvert\mathcal{N}\right\rvert}$. The problem is formulated as follows,
$\min_{z\geq 0}\quad\mathbb{E}\left[c^{\top}z+V(z,Y)\,\middle|\,X=x\right],$
where $V(z,Y)=\min_{f\geq 0}\left\{g^{\top}f+b^{\top}(Y+Wf-z)^{+}\right\}$.
This is a two-stage stochastic linear program, and its ERM problem can be converted into a large linear program by adding the second-stage decision variables for each scenario. Following the graph generation settings in Bertsimas and Koduri (2021), we set $\left\lvert\mathcal{N}\right\rvert=5$, $\left\lvert\mathcal{A}\right\rvert=10$, $c=(3,\dots,3)$, $g=(1,\dots,1)$, and $b=(5,\dots,5)$. To evaluate PADR’s performance under more complex conditions, we employed the following demand model:
$\overline{Y}^{\prime}(X)=k\alpha_{1}\sin(\alpha_{2}X_{1})+k\max\{\alpha_{3}X_{2},\alpha_{4}X_{2}\}+\alpha_{5},$ (19)
where the sine function represents the seasonal dependence of the demand on $X_{1}$, and the max-affine function represents the nonlinear trend with respect to $X_{2}$. We set $k=1$ unless otherwise specified. We randomly generate parameters $\{\alpha_{i}\}_{i=1}^{5}$ and the node-arc matrix $W$ for each trial, and we report performance of methods over 20 trials for each experimental setting. The sof method from Kallus and Mao (2023) is not applied for this two-stage SP, as it would require a new development of second-order perturbation analysis of stochastic optimization, which is beyond the scope of this paper.
4.3.1 Varying feature dimensions.
We first study the effect of the feature dimension $p$ on the performance of PADR under sparse and dense demand models similar to the discussion in Section 4.1.2. The results are reported in Figure 1p. Similar to the results in Section 4.1.2 for the Newsvendor problem, with sparse demand model, PADR performs well in low-dimensional cases but deteriorates as $p$ increases which may necessitate the proper parameter regularization. In contrast, forest-based prescriptive methods are robust to redundant features. With dense demand model, PADR and PO-PA have the best performance, with the former one slightly outperforming when the feature dimension increases to 50. Despite the increased complexity and the challenge in accurately approximating the true model, PADR continues to demonstrate superior performance for instances with dense demand model, high nonlinearity, and sparse demand models with small feature dimensions.
[…]
(o) Sparse demand model
(p) Dense demand model
Product placement problem with varying feature dimensions
4.3.2 Nonlinearity in demand model.
We also study the effect of nonlinearity in the demand model by varying the coefficient $k$. Notice that the nonlinear structure of the demand model here is different from the one in Section 4.1 with the sine component function. The results for $p=20$ are provided in Figure 1r. The increasing gap in Figure 1q indicates that except for PADR and PO-PA, the performance of all the other methods deteriorates as the nonlinearity of the demand model increases. Figure 1r indicates that PADR is slightly better than PO-PA and outperforms the other methods, and the relative gap among them remains approximately stable as $k$ increases.
[…]
Product placement problem: varying nonlinearity level $k$ (dense model, $p=20$) In (b), with the cost of the cart scaled to 1 and the cost of SIMOPT set to 0, we plot the relative cost of other methods compared to SIMOPT.
@@ 5 Conclusion
We propose the PADR-based ERM method together with the ESMM algorithm to solve a broad class of SP problems with covariate information. For our PADR-based ERM framework, we provide the approximation bound of the PA function class and further establish the nonasymptotic consistency result for unconstrained problems. To solve the ERM problem with a composite DC structure, we develop the ESMM algorithm by combining enhanced $\varepsilon$-active surrogation with sequential sampling. We provide the asymptotic convergence of the ESMM algorithm to composite strong d-stationarity, and the nonasymptotic convergence to d-stationarity via the error-bound analysis. Numerical experiments indicate the advantages of our method over PO methods and prominent DR-based methods for various instances regarding the test cost, computational time, and robustness to nonlinearity of the underlying dependency. In conclusion, the PADR is a favorable hypothesis class for DR-based ERM methods, supported by theoretical consistency, algorithmic convergence, and computational efficiency for solving SP with covariate information. Extensions of the PADR-based ERM framework for more complicated SP problems are deserved to be explored in the future, such as problems with chance constraints or endogenous random variables, which necessitate a further investigation of model construction, consistency analysis, and data-driven algorithmic development for problems with more complex composite structures.
@@ 6 Proofs of Theoretical Consistency Results in Section 2
Proof 6.1
@@
[…]
@@
Let $\Phi$ be a class of functions mapping from $\mathcal{U}$ to $\mathbb{R}$ and $\Xi_{N}=\{u^{s}\}_{s=1}^{N}\subseteq\mathcal{U}$ be a sample set of size $N$. Then the empirical Rademacher complexity of $\Phi$ with respect to $\Xi_{N}$ is defined as
^$\widehat{\mathcal{R}}_{\Xi_{N}}(\Phi)\coloneqq\mathbb{E}_{\boldsymbol{\sigma}}\left[\sup_{\phi\in\Phi}\frac{1}{N}\sum_{s=1}^{N}\sigma_{s}\phi(u^{s})\right],$
@@
$\mathcal{R}_{N}(\Phi)\coloneqq\mathbb{E}_{\Xi_{N}}[\widehat{\mathcal{R}}_{\Xi_{N}}(\Phi)].$
[…]
@@
Proof 6.4
[…]
^$\widehat{\mathcal{R}}_{\Xi_{N}}(\widehat{\Phi})\leq L_{\widehat{\Phi}}\sqrt{q}\epsilon+M_{\widehat{\Phi}}\sqrt{\frac{2q\log(2\mu/\epsilon)}{N}}.$
By setting $\epsilon=2\mu/N$ and taking expectation over $\Xi_{N}$, we obtain the desired upper bound.∎
@@
[…]
Let $\Phi$ be a function class mapping from $\mathcal{U}$ to $[-M,M]$. Then for any $\delta>0$, with probability at least $1-\delta$ over the draw of an independently and identically distributed (i.i.d.) sample $\Xi_{N}$ of size $N$, the following holds true for all $\phi\in\Phi$:
@@
Proof 6.6
Proof of Proposition 2.2. We first quantify the Rademacher complexity of the composite function class $\Phi_{{\mathcal{H}}_{\mu}^{K}}\coloneqq\left\{\phi:\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}\mid\phi(x,y)=\varphi({f(x)},{y}),\,f\in{\mathcal{H}}_{\mu}^{K}\right\}$. Notice that any function in $\Phi_{{\mathcal{H}}_{\mu}^{K}}$ is parameterized by some $\theta\in[-\mu,\mu]^{2dK(p+1)}$. For any $(x,y)\in\mathcal{X}\times\mathcal{Y}$ and $\theta_{1},\theta_{2}\in[-\mu,\mu]^{2dK(p+1)}$,
$\left\lvert\varphi(f(x,\theta_{1}),y)-\varphi(f(x,\theta_{2}),y)\right\rvert\leq L_{\varphi}\left\lVert f(x,\theta_{1})-f(x,\theta_{2})\right\rVert\leq 2L_{\varphi}\sqrt{p\bar{X}^{2}+1}\left\lVert\theta_{1}-\theta_{2}\right\rVert,$
@@
$\left\lvert\varphi(f(x),y)\right\rvert\leq\left\lvert\varphi(z_{0},y)\right\rvert+\left\lvert\varphi(f(x),y)-\varphi(z_{0},y)\right\rvert\leq C_{\varphi}+L_{\varphi}\left\lVert f(x)-z_{0}\right\rVert\leq M_{\mu}.$ (21)
@@
$\max_{f\in{\mathcal{H}}_{\mu}^{K}}\left\lvert R(f)-\widehat{R}_{\Xi}(f)\right\rvert\leq 2\mathcal{R}_{n}(\Phi_{{\mathcal{H}}_{\mu}^{K}})+\sqrt{2}M_{\mu}\sqrt{\frac{\log(2/\delta)}{n}}.$ (23)
@@
$=\mathbb{E}_{X,Y}\left[\varphi({f_{app}(X)},{Y})-\varphi({f^{*}(X)},{Y})\right]$
@@
$R(f^{*}_{\Xi,{\mathcal{H}}_{\mu}^{K}})-R(f^{*}_{{\mathcal{H}}_{\mu}^{K}})$ $=R(f^{*}_{\Xi,{\mathcal{H}}_{\mu}^{K}})-\widehat{R}_{\Xi}(f^{*}_{\Xi,{\mathcal{H}}_{\mu}^{K}})+\widehat{R}_{\Xi}(f^{*}_{\Xi,{\mathcal{H}}_{\mu}^{K}})-R(f^{*}_{{\mathcal{H}}_{\mu}^{K}})$
$\leq R(f^{*}_{\Xi,{\mathcal{H}}_{\mu}^{K}})-\widehat{R}_{\Xi}(f^{*}_{\Xi,{\mathcal{H}}_{\mu}^{K}})+\widehat{R}_{\Xi}(f^{*}_{{\mathcal{H}}_{\mu}^{K}})-R(f^{*}_{{\mathcal{H}}_{\mu}^{K}})$
$\leq 2\left(\max_{f\in{\mathcal{H}}_{\mu}^{K}}\,|R(f)-\widehat{R}_{\Xi}(f)|\right).$
@@
$R(f^{*}_{\Xi,{\mathcal{H}}_{\mu}^{K}})-R(f^{*}_{{\mathcal{H}}_{\mu}^{K}})\leq\frac{16\mu L_{\varphi}\sqrt{2dK(p+1)(p\bar{X}^{2}+1)}}{n}+4M_{\mu}\sqrt{\frac{2dK(p+1)\log n}{n}}+2\sqrt{2}M_{\mu}\sqrt{\frac{\log(2/\delta)}{n}}.$
@@
7 Surrogation Construction of $F(\cdot,\xi)$ for Case 1 to Case 3
We provide the construction of surrogation of $F(\cdot,\xi)\coloneqq\varphi(f(\cdot,\xi),\xi)$ for Case 1 to Case 3 in the proof of Proposition 3.2. In order to verify Assumption (B4), we prove that the minimum surrogate function $\widehat{F}^{\varepsilon}_{\min}(\cdot,\xi;\bar{\theta})$ at a reference point $\bar{\theta}$ satisfies the following condition for any $\xi\in\Xi$:
  1. (P4)